OpenStudy (anonymous):

A rectangular field has a perimeter of 100m. The length of the field is 10 m more than the breadth. find the area of the field. 1- 600sq.m 2- 2150sq.m 3- 2375sq.m 4- 2560sq.m

OpenStudy (butterflydreamer):

Let breadth = x , length = 10 + x Perimeter is the total lengths of all the sides added up basically :) In a rectangle we have 2 lengths and 2 breadths.  Therefore, Perimeter = 2*length + 2*breadth We know perimeter = 100 , length = 10 + x , breadth = x Sub it in and we get: 100 = 2 ( 10 + x) + 2x